The sum of two digits of a number is 11.
If the digits are reversed the number is 63 more than the original number.
What is the original number?

Answer by rfadrogane(214)

The sum of two digits of a number is 11.
If the digits are reversed the number is 63 more than the original number.
What is the original number?
Let: N = the original number.
N = 10t+u
t=the ten's digit
u-unit digit
Sol'n:
t+u=11 ----(1), the sum of two digits of a number is 11.
10u+t=(10t+u)+63, when the digits are reversed the number is 63 more than the original number.
9t-9u=-63, when divide by 9
t-u = -7----(2)
Add (1) and (2).
2t=4
t=2 and for u:
u=11-2
u=9
thus,
N=10(2)+9
N=29 ----answer
Check:
92-29=63
^_^